### v1

@@ Abstract
Surgical navigation based on multimodal image registration has played a significant role in providing intraoperative guidance to surgeons by showing the relative position of the target area to critical anatomical structures during surgery. However, due to the differences between multimodal images and intraoperative image deformation caused by tissue displacement and removal during the surgery, effective registration of preoperative and intraoperative multimodal images faces significant challenges. To address the multimodal image registration challenges in Learn2Reg 2024, an unsupervised multimodal medical image registration method based on multilevel correlation balanced optimization (MCBO) is designed to solve these problems. First, the features of each modality are extracted based on the modality independent neighborhood descriptor, and the multimodal images is mapped to the feature space. Second, a multilevel pyramidal fusion optimization mechanism is designed to achieve global optimization and local detail complementation of the deformation field through dense correlation analysis and weight-balanced coupled convex optimization for input features at different scales. For preoperative medical images in different modalities, the alignment and stacking of valid information between different modalities is achieved by the maximum fusion between deformation fields. Our method focuses on the ReMIND2Reg task in Learn2Reg 2024, and to verify the generality of the method, we also tested it on the COMULIS3DCLEM task. Based on the results, our method achieved second place in the validation of both two tasks.
@@ 1 Introduction
Medical image registration has been an important topic in the field of medical image analysis, and many significant methods [1, 2, 3, 5] have driven the development of medical image registration tasks. Deep learning-based medical image registration methods [4] usually involve long and complex learning processes, and often struggle to achieve accurate estimation for multimodal, large-deformation data and general usability for extensive tasks. The sub-challenge of Learn2Reg 2024, ReMIND2Reg, is a multimodal medical image registration task oriented to preoperative ultrasound and intraoperative MRI, which is characterized as unlabeled, large deformation, and low feature distinctness. Aiming at the above characteristics, inspired by [6, 7], an unsupervised multimodal medical image registration method based on multilevel correlation balanced optimization (MCBO) has been proposed, which can quickly achieve the effective registration of multimodal medical images by only a small number of learning and optimization procedures.
@@ 2 Methodology
The proposed MCBO method is based on convexAdam [8] with a series of improvements, which include (1) Introducing a weight balancing term on the coupled convex optimization to achieve smoother deformation optimization. (2) A multilevel pyramidal fusion optimization mechanism is designed to achieve the refinement of the dense deformation field by fusing the optimization results of different scales. (3) For the multimodal preoperative medical images in the task, the alignment and stacking of valid information between different modalities is achieved through the maximum fusion of deformation fields.
The overall flow of the proposed MCBO method is shown in Fig. 1. The moving image and the fixed image are inputted and then the modal-independent features of the images are first obtained by Mind-SSC Feature Extractor [7]. The Mind-SSC feature extractor exploits the self-similarity of partial area in the image to extract the unique structural information of the local neighborhood, which results in a highly consistent structural representation across modalities.
@@
The acquired features are subsequently downsampled by average pooling operations with different times $n$ ($n=1,2,4,6$) to obtain leveled features $Fi_{1}$ to $Fi_{4}$ as inputs for multilevel pyramidal fusion optimization mechanism. For the input features of levels 2 to 4, the feature matrix is directly passed through the dense correlation layer, the weight-balanced coupled convex optimization layer, and the inverse-consistency constraint layer to obtain the initial deformation fields of each level. For the first level where the input features are of the original scale, a pyramid splitting optimization strategy is used to reduce the amount of computation during the optimization procedure and at the same time to enhance the ability to align the low-contrast features. Specifically, the original feature is first divided into eight parts by equally splitting along the H, W, D dimensions of the input feature respectively, and the same three-step process as the other levels is performed for each part of the feature separately, followed by splicing the feature in the original dimensions to get the initial deformation field of the first level. To ensure the smoothness of the deformation field after splicing, an additional adam instance optimization [6] operation is performed on the output of the first level. Finally, the deformation fields at each level are summed up by multilevel weighted fusion using different weights, where the sum of the weights is 1, to obtain the final deformation field.

### v2

@@ Abstract
Surgical navigation based on multimodal image registration has played a significant role in providing intraoperative guidance to surgeons by showing the relative position of the target area to critical anatomical structures during surgery. However, due to the differences between multimodal images and intraoperative image deformation caused by tissue displacement and removal during surgery, effective registration of preoperative and intraoperative multimodal images faces significant challenges. To address the multimodal image registration challenges in Learn2Reg 2024, an unsupervised multimodal medical image registration method based on multilevel correlation balanced optimization (MCBO) is designed to solve these problems. First, the features of each modality are extracted based on the modality independent neighborhood descriptor, and the multimodal images are mapped to the feature space. Second, a multilevel pyramidal fusion optimization mechanism is designed to achieve global optimization and local detail complementation of the deformation field through dense correlation analysis and weight-balanced coupled convex optimization for input features at different scales. For preoperative medical images in different modalities, the alignment and stacking of valid information between different modalities is achieved by the maximum fusion between deformation fields. Our method focuses on the ReMIND2Reg task in Learn2Reg 2024, and to verify the generality of the method, we also tested it on the COMULIS3DCLEM task. Based on the results, our method achieved second place in the validation of both two tasks. The code is available at https://github.com/wjiazheng/MCBO.
@@ 1 Introduction
Medical image registration has been an important topic in the field of medical image analysis, and many significant methods [1, 2, 3, 5] have driven the development of medical image registration tasks. Deep learning-based medical image registration methods [4] generally involve long and complex learning processes, and often struggle to achieve accurate estimation for multimodal, large-deformation data and general usability for extensive tasks. The Learn2Reg 2024 subchallenge, ReMIND2Reg, is a multimodal medical image registration task oriented to preoperative ultrasound and intraoperative magnetic resonance imaging (MRI), which is characterized as unlabeled, large deformation, and low feature distinctness. Aiming at the above characteristics, inspired by [6, 7], an unsupervised multimodal medical image registration method based on multilevel correlation balanced optimization (MCBO) has been proposed, which can quickly achieve effective multimodal medical image registration using only a small number of learning and optimization procedures.
@@ 2 Methodology
The proposed MCBO method is based on convexAdam [8] with a series of improvements, which include (1) Introducing a weight-balancing term on coupled convex optimization to achieve smoother deformation optimization. (2) A multilevel pyramidal fusion optimization mechanism is designed to achieve refinement of the dense deformation field by fusing the optimization results of different scales. (3) For multimodal preoperative medical images in the task, the alignment and stacking of valid information between different modalities is achieved through the maximum fusion of deformation fields.
The overall flow of the proposed MCBO method is shown in Fig. 1. The moving image and the fixed image are inputted and then the modal-independent features of the images are first obtained by Mind-SSC Feature Extractor [7]. The Mind-SSC feature extractor exploits the self-similarity of the partial area in the image to extract the unique structural information of the local neighborhood, which results in a highly consistent structural representation across modalities.
@@
The acquired features are subsequently downsampled by average pooling operations at different times $n$ ($n=1,2,4,6$) to obtain leveled features $Fi_{1}$ to $Fi_{4}$ as input for the multilevel pyramidal fusion optimization mechanism. For the input features of levels 2 to 4, the feature matrix is directly passed through the dense correlation layer, the weight-balanced coupled convex optimization layer, and the inverse-consistency constraint layer to obtain the initial deformation fields of each level. For the first level where the input features are of the original scale, a pyramid splitting optimization strategy is used to reduce the amount of computation during the optimization procedure and, at the same time, to enhance the ability to align the low-contrast features. Specifically, the original feature is first divided into eight parts by equally splitting along the H, W, D dimensions of the input feature respectively, and the same three-step process as the other levels is performed for each part of the feature separately, followed by splicing the feature in the original dimensions to get the initial deformation field of the first level. To ensure smoothness of the deformation field after splicing, an additional adam instance optimization [6] operation is performed on the output of the first level. Finally, the deformation fields at each level are summed up by multilevel weighted fusion using different weights, where the sum of the weights is 1, to obtain the final deformation field.
@@ 3 Experiments and Results
Table 2: Results of ReMIND2Reg sub-challenge in Learn2Reg 2024.
TRE(mm)
Initial 3.727 $\pm$ 0.714
ConvexAdam-Rigid 2.773 $\pm$ 1.273
NiftyReg 2.751 $\pm$ 1.333
ours(without fusion) 2.453 $\pm$ 0.613
ours(with fusion) 2.224 $\pm$ 0.639